Determine the surface area generated by revolving the arc 3y= 4x from x=0 to x=3 about the x-axis.


1
Expert's answer
2021-05-19T08:04:52-0400

3y = 4x

Surface Area generated by revolving the arc 3y= 4x from x=0 to x=3 about the x-axis is given by


Area = ∫03 2 "\pi" * y * "\\sqrt{\\smash[b]{1 + (dy\/dx)^2\ufeff }}" dx ..............Equation(1)


dy/dx = 4/3 = 1.33


Area = ∫03 2 "\pi" * 1.33x * "\sqrt{1 + 1.33^2}" dx


Area = ∫03 2 "\pi"* 1.33x * 1.664 dx

Area = 13.89 ∫03 x dx

Area = 13.89 * ( x2 / 2) 03

Area = 13.89 * 4.5


Area = 62.54 square units